OpenStudy (anonymous):

The x-coordinate of the point of intersection of the graphs of -2x+y=10 and 4x+y=4 is x=

hero (hero):

-2x+y=10 4x+y=4 y = 10 + 2x y = 4-4x y = y 10 + 2x = 4 - 4x 6x = -6 x = -6/6 x = -1

OpenStudy (anonymous):